OpenStudy (anonymous):

Convert the equation 5x^2+5y^2=3y from a rectangular equation to a polar equation

OpenStudy (earthcitizen):

by rectangular do you mean a+jb

OpenStudy (anonymous):

complete the square

OpenStudy (earthcitizen):

\[r<\theta \] ?

OpenStudy (amistre64):

r = sqrt(x^2+y^2) stuff

OpenStudy (amistre64):

x = r cos(..) y = r sin(...)

OpenStudy (amistre64):

5x^2+5y^2-3y=0 5(x^2) +5(y^2-3/5 y) = 0 5(x^2) +5(y^2-3/5 y +(9/25)) = 5(9/25) 5(x^2) +5(y-3/25)^2 = 5(9/25) x^2 + (y-3/25)^2 = 9/25

OpenStudy (amistre64):

looks like r = 3/5 in this case

OpenStudy (amistre64):

\[(\frac{3}{5},\frac{3}{5}sin(t))\] maybe?